. Express each number as a product of its prime factors: 
 
 7429

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

By Taking the LCM of 7429, we will get the product of its prime factor.

Hence,
7429 = 17 × 19 × 23 × 1 = 17 × 19 × 23

Thanks